# From Flat to Feeling: A Feasibility and Impact Study on Dynamic Facial Emotions in AI-Generated Avatars

### v1

@@ Abstract
Dynamic facial emotion is essential for believable AI-generated avatars; however, most systems remain visually inert, limiting their utility in high-stakes simulations such as virtual training for investigative interviews with abused children. We introduce and evaluate a real-time architecture fusing Unreal Engine 5 MetaHuman rendering with NVIDIA Omniverse Audio2Face to translate vocal prosody into high-fidelity facial expressions on photorealistic child avatars. We implemented a distributed two-PC setup that decouples language processing and speech synthesis from GPU-intensive rendering, designed to support low-latency interaction in desktop and VR environments. A between-subjects study ($N=70$) using audio+visual and visual-only conditions assessed perceptual impacts as participants rated emotional clarity, facial realism, and empathy for two avatars expressing joy, sadness, and anger.
Results demonstrate that avatars could express emotions recognizably, with sadness and joy achieving high identification rates. However, anger recognition significantly dropped without audio, highlighting the importance of congruent vocal cues for high-arousal emotions. Interestingly, removing audio boosted perceived facial realism, suggesting that audiovisual desynchrony remains a key design challenge. These findings confirm the technical feasibility of generating emotionally expressive avatars and provide guidance for improving non-verbal communication in sensitive training simulations.
@@ 1 Introduction
Figure 1: Example frames of the two avatars used in the user study: "Emory" (male-presenting) and "Amelia" (female-presenting). The top row displays their neutral facial expressions. The bottom rows show expressions of joy (left) and sadness (right) for each avatar, as used in the study.
@@
Addressing these limitations requires a focused research endeavor into developing AI avatar systems capable of rich, dynamic facial emotion expression with real-time responsiveness. This research directly tackles this objective by introducing and assessing a novel, distributed architecture engineered for high-fidelity emotional expressiveness. This system represents a significant advancement by integrating the sophisticated rendering capabilities of Unreal Engine 5 (UE5) MetaHuman software [4] with the intelligent, audio-driven facial animation provided by NVIDIA’s Omniverse Audio2Face [5]. The system’s distributed architecture employs two synchronized computational nodes to effectively balance processing demands—decoupling language understanding and speech generation from GPU-intensive rendering—thereby enabling low-latency interaction. Figure 1 displays representative screenshots of the two photorealistic MetaHuman avatars, "Amelia" and "Emory," utilized in this study.
In summary, our main contributions are as follows:
  • Designing and implementing a modular pipeline for generating emotionally expressive avatars, emphasizing real-time facial animation using NVIDIA Audio2Face and Unreal Engine MetaHuman.
  • Conducting a user study ($N=70$) comparing audio+visual and visual-only conditions, highlighting trade-offs between cross-modal interference and perceived realism.
  • Demonstrating that high-arousal emotions (e.g., anger) rely heavily on audio-visual congruence, while low-arousal emotions (e.g., sadness) achieve robust recognition through visual cues alone.
  • Quantifying how facial structure (angular vs. soft) shapes emotion perception, guiding context-specific avatar design.
  • Identifying key challenges in avatar design and implementation, such as synchronization, the uncanny valley, and expression subtlety, through qualitative feedback analysis.
@@ 3.1 System Design and Implementation
To overcome the limitations of previous systems [3] for emotional expression in AI avatars, this research introduces a novel architecture that integrates Unreal Engine 5 Metahuman [4] for high-fidelity facial rendering with NVIDIA’s Omniverse Audio2Face [5] for generating emotionally expressive facial animations from audio input. Below, we provide a brief overview of these two core platforms:
@@ 3.2 User Study
A total of 70 participants were recruited via the Prolific platform and randomly assigned to one of two experimental conditions ($N=35$ per condition). The sample exhibited diversity in gender, age, academic/professional background, and familiarity with virtual environments.
@@
Participants viewed a series of six high-definition video clips, the presentation order of which was counterbalanced to avoid order effects. In these clips, the avatars expressed three emotional states: sadness, anger, and joy (see Figure 3). These specific emotions were selected as they represent a range of valence and arousal levels and are relevant emotions that children might express in sensitive interview contexts. In the audio+visual condition, speech was synthesized using OpenAI’s TTS system to ensure consistent vocal tone and emotionally rich delivery. Since child or adolescent voices were not available, we used a young adult female voice for both avatars. This choice was supported by findings from our prior studies [33], which showed that female voices can be effectively used to represent child characters in training scenarios, regardless of gender.
@@
To manage participant expectations and establish a perceptual anchor, we began with a deliberately low-fidelity stimulus, following evidence that anchoring effects require the presence of both semantic and numerical cues [34]. The inclusion of this initial low-fidelity stimulus was intended to provide a baseline or reference point for participants, potentially making their subsequent ratings of the higher-fidelity main stimuli more calibrated and less susceptible to being skewed by initial unfamiliarity or overly high expectations for avatar realism. By first exposing them to a clearly less sophisticated avatar, we aimed to help participants better contextualize and evaluate the nuances of the emotionally expressive avatars that formed the core of the study.
After viewing the main stimuli, participants completed a questionnaire comprising three main sections (see Table 1). Most items were rated on a five-point Likert scale and assessed aspects such as emotional clarity (how well the intended emotion was recognized), realism of facial expressions, and perceived emotional connection or empathy felt toward the avatars. The final section included an open-ended question inviting participants to reflect on what stood out most about the avatars.
Table 1: Structure and Content of the User Study Questionnaire
@@ 4 Results and Analysis
This section presents the findings of the user study, analyzing how participants perceived emotional expressions across different conditions and avatars, based on both quantitative ratings and qualitative feedback.
To evaluate perceptual differences across experimental conditions and avatars, statistical comparisons were performed using independent-samples $t$-tests or Welch’s $t$-tests. The specific test was chosen based on the results of Levene’s test for equality of variances. This approach ensures statistically robust comparisons while appropriately accounting for potential heteroscedasticity, as recommended by Field et al. [35]. The outcomes of these comparisons are detailed in Table 2 (for audio+visual vs. visual-only conditions), Table 3 (for general questionnaire items), and Table 4 (for comparisons between the two avatars).
Descriptive statistics, specifically means ($M$) and standard deviations ($SD$), were calculated for each variable under investigation. These descriptive data are graphically visualized in Figure 4. For enhanced clarity and direct comparison, all aforementioned tables (Table 2, Table 3, and Table 4) also present these mean values alongside the corresponding inferential test statistics, such as $t$-statistics and $p$-values.
4.1 Comparison of Audio and Visual Versus Visual-Only Conditions
@@
Figure 4: Mean ratings ($\pm$±SD) of emotional expressiveness across five emotions (Angry, Sad, Joy, Fear, Disgust) for the avatars Emory and Amelia, under both audio+visual and visual-only conditions. Each video clip is labeled on the left, with color-coded bars distinguishing the experimental conditions. Green-highlighted boxes indicate intended emotions whose recognition was especially important for the study.
@@
Variable t-statistic p-value
Mean
Audio+Visual
Mean
Visual-Only
Facial exp. 2.843 0.006 2.800 3.429
Angry -2.405 0.019 2.571 1.829
Sad -0.923 0.360 3.629 3.343
Joy 1.579 0.119 1.171 1.429
Fear 1.318 0.192 1.857 2.229
Emory_Sad Disgust 0.000 1.000 1.971 1.971
Facial exp. 2.239 0.028 2.914 3.514
Angry -10.729 0.000 3.829 1.429
Sad -4.122 0.000 3.257 2.000
Joy 5.559 0.000 1.086 2.143
Fear -0.762 0.448 1.914 1.714
Emory_Angry Disgust -4.306 0.000 2.743 1.486
Facial exp. 1.222 0.226 3.171 3.514
Angry 2.766 0.008 1.057 1.400
Sad 3.000 0.005 1.114 1.743
Joy -2.887 0.005 3.771 2.914
Fear 1.768 0.082 1.229 1.571
Emory_Joy Disgust 2.743 0.009 1.171 1.743
Facial exp. 2.836 0.006 3.343 4.000
Angry -0.803 0.425 2.571 2.314
Sad -0.125 0.901 4.114 4.086
Joy 1.226 0.224 1.086 1.200
Fear 0.751 0.455 2.086 2.314
Amelia_Sad Disgust -2.145 0.036 2.429 1.800
Facial exp. 0.351 0.726 3.543 3.629
Angry -2.988 0.004 3.114 2.171
Sad -4.172 0.000 3.714 2.371
Joy 2.928 0.005 1.143 1.829
Fear -0.543 0.589 2.000 1.857
Amelia_Angry Disgust -0.560 0.577 2.114 1.943
Facial exp. 1.680 0.098 3.486 3.914
Angry 1.358 0.179 1.000 1.086
Sad 0.848 0.400 1.057 1.143
Joy 0.838 0.405 4.200 4.371
Fear 0.314 0.754 1.114 1.143
Amelia_Joy Disgust 1.577 0.119 1.057 1.229
The analysis of perceptual differences between the audio+visual and visual-only conditions is structured as follows:
  • Perceived Facial Expressions Realism: As shown in Table 2, a consistent pattern emerges in participants’ ratings of facial expression realism across conditions. Participants rated the visual realism of facial expressions significantly higher in the visual-only condition across all six stimulus clips. Mean scores for "Facial expressions Realism" uniformly increased when auditory input was removed. Statistically significant differences were observed in three of the six clips: Emory-Sad ($p=0.006$), Emory-Angry ($p=0.028$), and Amelia-Sad ($p=0.006$). This phenomenon likely arises from audiovisual interference; the presence of synthetic speech may divide attentional resources between auditory and visual channels, thereby amplifying minor artifacts such as lip desynchronization. In the visual-only condition, the unimodal evaluation eliminates such cross-modal incongruities, allowing participants to assess facial animations in isolation, which consequently enhances perceived realism.
  • Recognition of Sadness and Joy: Findings presented in Table 2 and Figure 4 indicate the emotions of sadness and joy were consistently recognized across both experimental conditions, with both avatars achieving near-ceiling recognition scores for these emotions. For Amelia, sadness recognition remained stable between the audio+visual ($M=4.11$) and visual-only ($M=4.09$) conditions. Her joy recognition scores showed a slight, non-significant increase in the visual-only condition ($M=4.37$) compared to the audio+visual condition ($M=4.20,p=0.41$). In contrast, Emory’s sadness recognition showed minimal variation between the audio+visual ($M=3.63$) and visual-only ($M=3.34$) conditions. However, his joy recognition declined sharply and significantly in the visual-only condition, dropping from $M=3.77$ to $M=2.91$ ($p=0.005$).
    Sadness recognition remained stable across conditions, consistent with its reliance on visual cues like brow depression and lip-corner droop (FACS) [36]. Joy recognition, however, diverged by avatar: Amelia’s scores were robust in both conditions, likely due to clear Duchenne markers (e.g., cheek elevation), while Emory’s joy recognition dropped 22% in visual-only condition. This suggests participants relied on auditory prosody (e.g., rising pitch) to distinguish Emory’s joy from frustration; in the visual-only condition, his brow tension was misinterpreted as negative affect, underscoring the interplay between modality and avatar morphology. These findings confirm that sadness relies on visual fidelity, while joy benefits from audiovisual congruence, particularly through Duchenne smiles. Auditory cues minimally affected sadness but were critical for disambiguating Emory’s joy, aligning with prior work on facial action units (AUs) in low-arousal emotions [2]. The variance in joy recognition highlights modality’s context-dependent effects. Amelia’s visual clarity reduced cross-modal interference, improving joy perception without audio. Conversely, Emory’s angular morphology required vocal reinforcement to resolve ambiguity, exposing risks in incongruent multimodal design. This asymmetry challenges universal design rules, advocating instead for emotion-specific optimization: prioritize audiovisual synchronization for high-arousal emotions (e.g., joy) and visual precision for low-arousal ones (e.g., sadness). Such tailored approaches could mitigate mismatches between morphology and acoustic cues, advancing emotionally intelligent avatar design.
  • Recognition of Anger: Table 2 and Figure 4 show that anger recognition most strongly relied on audio cues for both avatars. Emory’s anger scores dropped from $M=3.83$ (audio+visual) to $M=1.43$ (visual-only), and Amelia’s from $M=3.11$ to $M=2.17$. These significant declines (Table 2; Emory: $p<0.001$, Amelia: $p=0.004$) accompanied frequent misclassifications of silent anger displays as mild sadness, with brow furrows seen as worry, not aggression. This aligns with appraisal theory and acoustic studies, indicating high-arousal emotions like anger depend on prosodic cues (e.g., vocal intensity, spectral noise). Without these, facial scowls alone were insufficient for disambiguation, stressing the need for multimodal congruence. Notably, even with audio, Amelia’s anger ($M_{anger}=3.11$) was perceptually overshadowed by sadness ($M_{sadness}=3.71$), unlike Emory’s anger-dominant profile ($M_{anger}=3.83$ vs. $M_{sadness}=3.26$), suggesting avatar morphology and voice design also modulate emotional salience.
    This perceptual asymmetry for anger likely stems from morpho-acoustic factors. Amelia’s softer features (e.g., jawline, subdued lip compression) may have reduced perceived aggression. Furthermore, her synthesized young female vocal profile lacked typical low-frequency anger cues, potentially biasing interpretations towards sorrow. These results reiterate the need for congruent multimodal signals for high-arousal emotions. Future work should investigate augmenting Amelia’s upper-face tension (e.g., sharper V-brows, AU4 activation) and integrating harsher prosody to enhance anger clarity without sacrificing fidelity.
  • Perceived realism and Empathy levels: The analysis presented in Table 3 explores two key dimensions: firstly, the influence of multimodal (audio and visual) cues on participants’ perception of realism, and secondly, their self-reported levels of empathy.
    In the audio+visual condition, auditory elements emerged as significant contributors to perceived realism. Specifically, voice tone received a mean rating of $4.0$, and dialogue content was rated at $M=3.86$. Conversely, aspects related to visual fidelity—namely facial expression realism and overall visual appearance—were rated higher in the visual-only condition ($M=3.86$ and $M=3.80$, respectively) compared to the audio+visual condition ($M=3.29$ and $M=3.40$, respectively). The difference in perceived facial expression realism between the two conditions was statistically significant ($t=2.21,p=0.03$). This suggests that the exclusion of auditory input may have reduced cross-modal interference, thereby enabling participants to focus more intently on, and assign higher ratings to, the visual fidelity of the avatars.
    When empathy levels were analyzed independently, no statistically significant difference was found between the visual-only condition ($M=3.06$) and the audio+visual condition ($M=2.83$; $t(68)=0.75,p=0.46$). While the means suggest a slight numerical increase in reported empathy in the visual-only condition, this difference did not reach statistical significance. One might speculate that, had a stronger effect been present, it could align with prior research suggesting that a unimodal visual focus might enhance attention to subtle facial expressions, potentially deepening emotional engagement. However, several factors in the current study could have influenced these results. For instance, in the audio+visual condition, the use of a young adult female voice for child avatars (a necessary choice due to the unavailability of suitable child or adolescent synthesized voices) may have created a perceptual disconnect for some participants. Furthermore, any instances where the synthesized voice did not fully align with the intended emotion, or where subtle audiovisual mismatches were perceived, could have potentially detracted from the emotional impact in the audio+visual condition. These factors might have counteracted any benefits of congruent multimodal cues for empathy in this specific experimental setup.

### v2

# Multimodal Integration Challenges in Emotionally Expressive Child Avatars for Training Applications

@@ Abstract
Dynamic facial emotion is essential for believable AI-generated avatars; however, most systems remain visually inert, limiting their utility in simulations such as virtual training for investigative interviews with abused children. We introduce and evaluate a real-time architecture fusing Unreal Engine 5 MetaHuman rendering with NVIDIA Omniverse Audio2Face to translate vocal prosody into high-fidelity facial expressions on photorealistic child avatars. Due to limitations in synthetic voice options, both avatars were voiced using a young adult female TTS model, selected from two different systems in an attempt to better match each character. This compromise introduces a confounding factor. Voice-age mismatches and prosodic or emotional differences may disrupt audiovisual alignment. We implemented a two-PC setup decoupling language and speech synthesis from GPU-intensive rendering, designed to support low-latency interaction in desktop and VR environments. A between-subjects study ($N=70$) using audio+visual and visual-only conditions assessed perceptual impacts as participants rated emotional clarity, facial realism, and empathy for two avatars expressing joy, sadness, and anger.
Results show that while avatars could express emotions recognizably, particularly sadness and joy, recognition of anger dropped markedly without audio, highlighting the role of vocal cues in conveying high-arousal states. Interestingly, silencing the clips improved perceived realism by removing mismatches between facial animation and voice, especially where age or emotional tone were incongruent. These findings highlight that perceived believability hinges on the interplay between audiovisual congruence and facial geometry: a mismatched voice can undermine even well-crafted expressions, while a congruent one can enhance weaker visuals. This trade-off presents an ongoing challenge for designing emotionally coherent avatars in sensitive training contexts.
@@ 1 Introduction
perceived believability hinges on the interplay between audiovisual congruence and facial geometry
Figure 1: Example frames of the two avatars used in the user study: "Emory" (male-presenting) and "Amelia" (female-presenting). The top row displays their neutral facial expressions. The bottom rows show expressions of joy (left) and sadness (right) for each avatar, as used in the study. For clarity, only two emotions are visualized here; the full set of expressions, including anger, is available in the supplementary videos provided in the project’s GitHub repository [3].
@@
Addressing these limitations requires a focused research endeavor into developing AI avatar systems capable of rich, dynamic facial emotion expression with real-time responsiveness. This research directly tackles this objective by introducing and assessing a novel, distributed architecture engineered for high-fidelity emotional expressiveness. This system represents a significant advancement by integrating the sophisticated rendering capabilities of Unreal Engine 5 (UE5) MetaHuman software [5] with the intelligent, audio-driven facial animation provided by NVIDIA’s Omniverse Audio2Face [6]. The system’s distributed architecture employs two synchronized computational nodes to effectively balance processing demands by decoupling language understanding and speech generation from GPU-intensive rendering, thereby enabling low-latency interaction. Figure 1 displays representative screenshots of the two photorealistic MetaHuman avatars, "Amelia" and "Emory," utilized in this study.
In summary, this study offers the following contributions:
  • Development of a modular, real-time pipeline integrating Unreal Engine MetaHuman and NVIDIA Audio2Face to generate emotionally expressive child avatars for interactive applications.
  • Empirical investigation ($N=70$) comparing audio+visual and visual-only conditions, revealing how mismatched modalities can reduce perceived realism, and in some cases, visual-only presentations may be preferred.
  • Identification of a perceptual asymmetry across emotional arousal levels: while low-arousal emotions (e.g., sadness, joy) were recognizable from facial cues alone, high-arousal emotions (e.g., anger) were more dependent on audiovisual congruence.
  • Observation that avatar facial morphology (e.g., soft vs. angular features) influences emotion interpretation, suggesting the need for morphology-emotion alignment in sensitive contexts such as empathy training.
  • Extraction of practical design lessons from participant feedback, including the impact of audiovisual desynchrony, the trade-off between expression clarity and naturalism, and the importance of age-appropriate voice synthesis.
@@ 3.1 System Design and Implementation
To overcome the limitations of previous systems [4] for emotional expression in AI avatars, this research introduces a novel architecture that integrates Unreal Engine 5 Metahuman [5] for high-fidelity facial rendering with NVIDIA’s Omniverse Audio2Face [6] for generating emotionally expressive facial animations from audio input. The source code and implementation details for the system described in this study are publicly available [3]. Below, we provide a brief overview of these two core platforms:
@@
    In this study, the “exaggerated” mode in Audio2Face was used to enhance the clarity and distinctiveness of facial expressions in short-duration clips, facilitating consistent emotion recognition among participants. This setting was chosen based on preliminary observations during development, as the standard and automatic modes often produced expressions that were too subtle to be reliably perceived within the limited duration of the stimuli.
@@ 3.2 User Study
A total of 77 participants were initially recruited via the Prolific platform. Although initial eligibility was filtered using Prolific’s pre-screening tools (e.g., age, language fluency, and device requirements), an additional quality control step was applied during data cleaning. Based on clearly invalid or inconsistent responses (e.g., repeated ratings across stimuli or contradictory answers), 7 participants were excluded from further analysis. The final sample included 70 participants whose responses met the criteria for attentiveness and internal consistency. Participants were randomly assigned to one of two experimental conditions ($N=35$ per condition) using Prolific’s built-in randomization tool. To minimize order effects, we manually randomized the sequence of avatar stimuli within each condition, varying both avatar identity and emotional expression. All participants in a given condition viewed the same randomized stimulus order. The sample exhibited diversity in gender, age, academic/professional background, and familiarity with virtual environments.
@@
Participants viewed a series of six high-definition video clips, the presentation order of which was counterbalanced to avoid order effects. In these clips, the avatars expressed three emotional states: sadness, anger, and joy (see Figure 3). These specific emotions were selected as they represent a range of valence and arousal levels and are relevant emotions that children might express in sensitive interview contexts. In the audio+visual condition, speech was synthesized using two different TTS systems: OpenAI’s TTS [^1] https://ttsopenai.com/ for the Emory avatar and NaturalReaders [^2] https://www.naturalreaders.com/ for the Amelia avatar. We evaluated several available text-to-speech options and selected these two voices based on their perceived suitability, as judged by two of the researchers. Although child-like voices were technically available, a young adult female voice was ultimately preferred for both avatars due to its better perceptual match with the visual and emotional characteristics of the characters. This choice was supported by findings from our prior studies [34], which showed that female voices can be effectively used to represent child characters in training scenarios, regardless of gender. Additionally, we found the child-like voices to be overly juvenile in tone and pitch, which created a mismatch with the avatars’ appearance and demeanor, both of which were designed to resemble pre-teens or young adolescents rather than very young children.
@@
Figure 4: Low-fidelity anchoring stimulus used at the beginning of the user study to calibrate participants’ expectations regarding avatar realism.
@@
To manage participant expectations and establish a perceptual anchor, we began with a deliberately low-fidelity stimulus, following evidence that anchoring effects require the presence of both semantic and numerical cues [35]. The inclusion of this initial low-fidelity stimulus was intended to provide a baseline or reference point for participants, potentially making their subsequent ratings of the higher-fidelity main stimuli more calibrated and less susceptible to being skewed by initial unfamiliarity or overly high expectations for avatar realism. An example frame from this anchoring stimulus is shown in Figure 4. By first exposing them to a clearly less sophisticated avatar, we aimed to help participants better contextualize and evaluate the nuances of the emotionally expressive avatars that formed the core of the study.
After viewing the main stimuli, participants completed a questionnaire comprising three main sections (see Table 1). items were rated on a five-point Likert scale. Emotion recognition and perceived facial realism were assessed for each individual video, while perceived realism across different contributing aspects (e.g., facial expressions, voice tone) and overall empathy toward the avatars were measured in a final, post-video section of the questionnaire. The final section included an open-ended question inviting participants to reflect on what stood out most about the avatars.
Table 1: Structure and content of the user study questionnaire, including measures of perceived realism, emotional expressiveness, and overall impressions. The full questionnaire is available in the project’s GitHub repository [3].
@@
3.3 Ethical Considerations
This study involved the evaluation of emotional expressiveness in synthetic avatars by adult participants via anonymous online surveys. All visual and auditory stimuli were fully computer-generated and did not include any real individuals or personal data. This study did not require formal ethical approval, as it involved no interaction with vulnerable groups, no collection of sensitive personal data, and no deception. Participants were recruited through the Prolific platform and were informed about the nature and duration of the study prior to participation. No formal consent form was presented beyond Prolific’s standard participation agreement. While the study featured simulated emotional expressions (e.g., sadness and anger) in photorealistic child avatars, no content warnings were provided prior to exposure. Some participants reported feeling uncomfortable or uneasy, particularly in response to expressions of sadness or anger. This is noted as a potential ethical consideration. Future implementations should consider including content advisories to reduce the risk of discomfort for sensitive viewers.
@@ 4 Results and Analysis
To comprehensively assess how participants interpreted the emotional expressions of the avatars, we conducted both quantitative and qualitative analyses.
4.1 Quantitative Analysis
The quantitative analysis in this study focused on how participants perceived and interpreted the emotional expressions of AI-generated child avatars across different experimental conditions. Specifically, we examined three main response categories:
  1. 1.
    Emotion recognition, defined as how accurately participants identified the intended emotional expression (anger, sadness, joy) after each video,
  2. 2.
    Perceived facial expression realism, referring to the extent to which facial animations appeared realistic in each video, and
  3. 3.
    Perceived realism across contributing aspects (e.g., facial expressions, visual appearance, voice tone, dialogue content), along with the empathic connection participants felt toward the avatars, assessed in a post-exposure questionnaire.
We first computed descriptive statistics (mean and standard deviation) for each category. To assess effects of condition, avatar identity, and emotion scenario on these outcomes, we employed cumulative link mixed-effects models (CLMMs) using the ordinal package in R. These models incorporated fixed effects for Condition (audio+visual vs. visual-only), Avatar (Amelia vs. Emory), and Scenario (anger, sadness, joy), along with their interaction terms. Random intercepts were included for participants and, where applicable, for emotion or video stimuli, to account for within-subject and item-level variability. Model fitting was performed using a logit link and adaptive Gauss–Hermite quadrature (nAGQ = 10) to ensure robust convergence and estimation accuracy.
Before examining each response category in detail, we begin by validating the robustness and goodness-of-fit of our statistical models. The results are then presented across the following dimensions: emotion recognition, perceived facial realism, and perceived realism/empathy. For completeness and cross-method validation, supplementary t-test analyses are reported in Appendix Table 8 and Table 9.
@@ 4.1 Quantitative Analysis
Figure 5: Mean ratings ($M\pm SD$) of emotional expressiveness across five emotions (Angry, Sad, Joy, Fear, Disgust) for the avatars Emory and Amelia, under both audio+visual and visual-only conditions. Bars represent group means; values next to each bar report the mean and standard deviation, while horizontal error bars indicate 95% confidence intervals (CI). Each video clip is labeled on the left, with color-coded bars distinguishing the experimental conditions. Green-highlighted boxes indicate target emotions critical for the study’s recognition task.
4.1.1 Statistical Validation and Model Robustness
To evaluate the reliability and robustness of the statistical modeling framework, we performed an extensive set of diagnostic procedures, model comparisons, and fit assessments across all three primary response categories. These included emotion recognition, perceived facial realism, and post-exposure ratings of realism contributors and empathic engagement.
For the emotion recognition analysis, model selection was guided by information-theoretic criteria, specifically the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC), along with likelihood ratio tests (LRTs) to assess violations of the proportional-odds (PO) assumption. As shown in Table 2, relaxing the Proportional Odds (PO) constraint significantly improved model fit ($p<0.001$), justifying the use of more flexible specifications in the CLMMs. To ensure sufficient power, an a priori simulation-based power analysis was conducted, confirming that a sample size of $N=70$ was adequate to detect medium-sized effects with a power of at least 0.80.
Similarly, Table 3 summarizes model comparisons for perceived facial expression realism. Again, significant improvements were observed when PO constraints were relaxed (e.g., LRT = 33.35, df = 15, $p=0.004$), indicating that ordinal responses were not evenly distributed across thresholds. The best-performing CLMMs, detailed in Table 4, achieved a conditional $R^{2}$ of 0.640, which indicates strong explanatory power when both fixed and random effects were considered. The marginal $R^{2}$ of 0.099 reflects the variance attributable to fixed effects alone. Residual diagnostics revealed an RMSE of 3.176 and residual standard deviation ($\sigma=3.238$), confirming acceptable levels of predictive error and model stability.
In addition to these analyses, we assessed model robustness for the third response category, which included perceived realism across contributing aspects (e.g., facial expressions, visual appearance, voice tone, dialogue content) and overall empathy toward the avatars. Internal consistency for the multi-item realism contributors scale was found to be marginal, with a Cronbach’s alpha of $\alpha=0.652$, suggesting moderate internal reliability. Furthermore, a logistic regression model predicting empathy responses yielded a Nagelkerke $R^{2}$ of just 0.011, indicating minimal variance explained. These findings suggest high individual variability and underscore the challenge of modeling subjective constructs like empathic connection using fixed predictors alone.
Across all three domains, model coefficients were exponentiated to derive interpretable odds ratios (OR) and visualized using forest plots. Where applicable, multiple comparisons were corrected using Holm’s method. To contextualize statistical findings, open-ended responses were triangulated with quantitative patterns, particularly regarding visual fidelity, emotional clarity, and audiovisual congruence. These effects were especially pronounced in high-arousal conditions, such as anger, where audiovisual alignment proved critical for accurate emotion recognition.
Taken together, these results validate the overall modeling approach and underscore its suitability for capturing the complex interplay between avatar design, experimental conditions, and emotional perception. The convergence of statistical rigor, diagnostic confirmation, and qualitative insight provides a robust foundation for interpreting user responses to emotionally expressive virtual avatars.
@@ 4.1.1 Statistical Validation and Model Robustness
Df logLik AIC LRT Pr(>Chisq) no.par
6 -2521.47 5076.95 30.57 0.00 17
Table 3: Model comparison for perceived facial expression realism. Relaxing the PO assumption improved model fit significantly across multiple model specifications.
Df logLik AIC LRT Pr(>Chi)
3.00 -382.84 937.69 9.34 0.02
6.00 -380.78 939.57 13.46 0.03
15.00 -370.84 937.69 33.34 0.004
Table 4: Fit statistics for the best-performing CLMM in the facial realism analysis. Conditional and marginal $R^{2}$ values indicate explained variance; RMSE and $\sigma$ reflect prediction error and residual variability.
AIC BIC R2_conditional R2_marginal RMSE $\sigma$
1035.65 1100.30 0.64 0.09 3.17 3.23
4.1.2 Emotion Recognition
To initiate our analysis of emotion recognition, we present the results of the ordinal mixed-effects model (Table 5), supported by two key visualizations: a forest plot illustrating the influence of model predictors (Figure 6) and a bar graph comparing perceived emotional expressiveness across experimental conditions and avatars identity (Figure 5). Together, these results offer a comprehensive overview of participants’ accuracy in identifying the target emotions conveyed by the avatars.
Figure 5 shows that sadness and joy were recognized with high accuracy across both conditions. Sadness recognition remained stable for both avatars (Amelia: $M=4.11$ vs. $M=4.09$; Emory: $M=3.63$ vs. $M=3.34$), consistent with its strong reliance on static visual cues such as brow depression (AU1+4) and lip-corner droop (AU15) as defined in FACS [36].
In contrast, recognition of joy varied by avatar identity and condition. Amelia’s joy remained consistently high, slightly improving in the visual-only condition ($M=4.37$ vs. $M=4.20$). Emory’s joy ratings, however, dropped 22% in visual-only condition ($M=3.77$ to $M=2.91$). This disparity may reflect both visual ambiguity and a potential methodological limitation, namely the use of two different TTS applications, which may have introduced variations in emotional prosody, and the use of adult female voices for both avatars, which might not have fully matched their child-like appearance. Emory’s voice, generated with OpenAI’s TTS, exhibited strong emotional expressiveness, particularly in terms of pitch variation and intensity. However, this high affective prosody may not have been fully matched by the corresponding facial expressions, which might have contributed to a perceived mismatch between voice and visuals. In contrast, Amelia’s voice, synthesized using NaturalReaders, seemed to align more closely with the visual and emotional tone of the avatar, although this could also be influenced by other factors such as facial morphology or expression timing. Nevertheless, we cannot attribute the entire effect to the voice alone, as the study also included visual-only conditions in which Amelia’s avatar consistently performed better in terms of emotion recognition and perceived realism. This suggests that additional factors, such as facial morphology and expression clarity, also played a substantial role. Amelia’s advantage may be partly explained by her softer facial features and the presence of clear Duchenne markers (e.g., AU6+12), which enhanced visual clarity even in the absence of audio. In contrast, Emory’s more angular facial structure may have contributed to perceptual ambiguity, with features such as brow tension being more easily misinterpreted as negative affect.